1. A machine produces metal rods used in an automobile suspension system. A random sample of 15 rods is selected, and the diameter is measured. The resulting data (in millimeters) are as follows: 11.44 11.39 11.35 11.38 11.45 11.4311.4211.44 11.34 11.39 11.46 11.36 11.44 11.49 11.41 a) Use a Normal Probability Plot to check the assumption of normality for rod diameter. b) Is it believable the mean rod diameter is less than 1 1.50 mm? Construct the appropriate 99% confidence interval
Answer #1

a)

Rank Proportion Z Value Data:
1 0.063 -1.534 11.34
2 0.125 -1.150 11.35
3 0.188 -0.887 11.36
4 0.250 -0.674 11.38
5 0.313 -0.489 11.39
6 0.375 -0.319 11.39
7 0.438 -0.157 11.41
8 0.500 0.000 11.42
9 0.563 0.157 11.43
10 0.625 0.319 11.44
11 0.688 0.489 11.44
12 0.750 0.674 11.44
13 0.813 0.887 11.45
14 0.875 1.150 11.46
15 0.938 1.534 11.49

Normal Probability Plot 11.5 11.45 11.4 11.35 ◆ Data: 132, 00 1,0000 2000 1.000 0.000 Z Value 1.000 2.000

the S shape of the graph in the normal probability plot indicates that distribution is approximately normal.

b)

mean=   11.41    [Excel function used -> AVERAGE]
sd=   0.05    [Excel function used -> STDEV]  
n=   15.00    [Excel function used -> COUNT]
alpha=   1%
critical value, t(a/2,n-1) = t(0.01/2,15-1) = 2.977

CI = mean +- t(a/2,n-1)*(sd/sqrt(n))      
lower = 11.4129 - 2.977*(0.045/sqrt(15)) = 11.38
upper = 11.4129 + 2.977*(0.045/sqrt(15)) = 11.45

i am 99% confident that estimated population mean rod diameter lie in this interval (11.38, 11.45). Since this interval has all values less than 11.5, i can say that, it is believable that mean rod diameter is less than 11.5.